OpenStudy (anonymous):

what would be the inverse function of ln(2+lnx) and it's domain? :)

OpenStudy (anonymous):

so do you know how to start finding the inverse function?

OpenStudy (anonymous):

yea i guess

OpenStudy (anonymous):

so how do you do it?

OpenStudy (anonymous):

we interchange x and y and solve the equation for y

OpenStudy (anonymous):

good and if you do that to your equation?

OpenStudy (anonymous):

that's the problem. I don't under stand how to actually do it for this function

OpenStudy (anonymous):

ahhh ok! well we start off with something like this correct: y=ln(2=lnx) as the original function

OpenStudy (anonymous):

*y = ln(2+lnx)

OpenStudy (anonymous):

yea but what after that? sorry but I'm kinda in a hurry :p

OpenStudy (anonymous):

I see well after switching we get x = ln(2+lny) and the part you probably are stuck on it how to switch over the ln (it's just the log scale for e) so we raise both sides by e. \[e^{x} = 2+\ln y\] think you can take it from here?

OpenStudy (anonymous):

what do you mean by raising both sides by e? raise it to the power e or take natural log on both sides?

OpenStudy (anonymous):

oh I'm sorry I meant that we take both sides and make it the power to e. so... \[e^{x} = e^{\ln (2 + \ln y)}\]

OpenStudy (anonymous):

and since lne =1 we get e^x=2+lny ?

OpenStudy (anonymous):

yes exactly

OpenStudy (anonymous):

sooo now we get y= e^(e^x-2) ? am I right? would this be the inverse?

OpenStudy (anonymous):

yup that's correct

OpenStudy (anonymous):

now to get the domain and range of inverse functions you simply reverse the two from the original function.

OpenStudy (anonymous):

domain of inverse would be range of original function so what would be the range of the original function?

OpenStudy (anonymous):

what is the range of just this function y= lnx ?

OpenStudy (anonymous):

0 to infinity. I think

OpenStudy (anonymous):

that's it's domain, the range is simply all real numbers (think about it's own inverse which is e^x: it's domain is all reals and it's range is [0,infinity)). So since lnx can be anything(aka all real numbers) 2+lnx can be all real numbers as well.

OpenStudy (anonymous):

oh sorry I mean (0,infinity) not [0,infinity)

OpenStudy (anonymous):

oh yes yes. sorry. So the the domain of our inverse function is simply the set of all real numbers?

OpenStudy (anonymous):

well that is the range of the original function correct so it must be and we can check it by looking at the inverse function itself, there are no sqrts or anything of the sort in the function and x is not in the denominator so it makes sense take it's the set of all real numbers.

OpenStudy (anonymous):

*that

OpenStudy (anonymous):

Oh yes i get it. Thank you SO VERY much! You've been great help.

OpenStudy (anonymous):

No problem so do you think you can solve for the range of the inverse or would you like my help?

OpenStudy (anonymous):

I think I can do it. And I just needed the domain anyway. But I'll do it just for the sake of it. Thanks again

OpenStudy (anonymous):

no problem good luck.

OpenStudy (anonymous):

:)